OpenStudy (anonymous):

select the two lines that are paralled y=3x y-1=3(x+2) -3x+4y=32 y=3/4x+8

OpenStudy (freckles):

first two have slope 3 second two have slope 3/4

OpenStudy (anonymous):

so there are paralled?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

1st and 2nd are parallel

OpenStudy (accessdenied):

-3x + 4y = 32 4y = 3x + 32 y = 3/4 x + 8 those lines wouldn't be parallel, they are the same line

OpenStudy (anonymous):

y=3x y-1=3(x+2) --> y = 3x + 7 Both have same slope m = 3 --> parallel (different y intercept)